# Prompt Engineer: Analyzing Skill Requirements in the AI Job Market

### v1

@@ Abstract
The rise of large language models (LLMs) has created a new job role: the Prompt Engineer. Despite growing interest in this position, we still do not fully understand what skills this new job role requires or how common these jobs are. We analyzed 20,662 job postings on LinkedIn, including 72 prompt engineer positions, to learn more about this emerging role. We found that prompt engineering is still rare (less than 0.5% of sampled job postings) but has a unique skill profile. Prompt engineers need AI knowledge (22.8%), prompt design skills (18.7%), good communication (21.9%), and creative problem-solving (15.8%) skills. These requirements significantly differ from those of established roles, such as data scientists and machine learning engineers, showing that prompt engineering is becoming its own profession. Our findings help job seekers, employers, and educational institutions in better understanding the emerging field of prompt engineering.
Keywords: Prompt engineering, large language models, LLM, AI workforce, skill analysis, job market analysis, text mining, natural language processing, human-AI interaction
@@ 1 Introduction
The rise of large language models (LLMs) has been transforming various industries (Chkirbene et al., 2024). AI enables advancements in automation, content generation, and AI-driven decision-making. However, the effectiveness of AI models heavily depends on how prompts are structured (Oppenlaender, 2023), leading to the emergence of prompt engineering as a specialized field. Companies are increasingly seeking to recruit prompt engineers in an effort to optimize and extract the full potential of LLMs. Despite the growing demand for this role, there remains no standardized definition of a prompt engineer, resulting in uncertainties regarding the scope, required skills, and industry relevance of this emerging job role.
Due to the timely relevance of LLM applications in industry, it is important to develop an understanding of the required knowledge and skills for this emerging job type. However, because of the novelty of AI technology, individuals who are interested in pursuing the role of the prompt engineer may face challenges understanding the qualifications required. Therefore, addressing this gap is beneficial to both job seekers and practitioners in industry as well as employers who want to identify suitable candidates. However, how can we systematically extract and analyze the skills, competencies, and industry relevance of this job role? In this paper, we employ a text mining approach to analyze job advertisements published on LinkedIn, aiming to uncover the key knowledge areas and skills required for prompt engineers. Understanding these trends provides insight into the broader adoption of LLMs and the pace of different industries in adapting to this evolving technology.
@@
  • RQ1:
    How prevalent are prompt engineer positions compared to more established job roles?
  • RQ2:
    What soft skills are most valued for prompt engineers, and how do these differ from other job roles?
  • RQ3:
    What technical skills define the prompt engineer role, and how do these hard skills create a distinct professional profile?
@@
The remainder of this paper is structured as follows. We review related literature on prompt engineering and the extraction of skills in Section 2. We then describe our approach to collecting and extracting information from job descriptions in Section 3. We then present our findings in Section 4, contrasting the skills of prompt engineers with fixe other related jobs. We discuss our findings in Section 5 and conclude in Section 6.
@@ 2.1 The Emergence of the Prompt Engineer
Artificial intelligence (AI) is changing the way workers perform their jobs and the skills they require (OECD, 2024; Brasse et al., 2024). Generative artificial intelligence has a strong impact on the future of work (Cazzaniga, 2024; Carroll, 2025) and higher education (Fan and Irawan, 2025). While concerns and anxiety over automation (Dinnen, 2022) are growing, AI also creates new professional opportunities (Elrawy and Wagdy, 2025; Kutela et al., 2023). There is a demand for a new digital workforce with a novel set of AI literacy (Long and Magerko, 2020) and digital skills (van Laar et al., 2017). One such novel job is the prompt engineer.
@@
Kutela et al. (2023) examined the nature, qualifications, and compensation of prompt engineers through an analysis of LinkedIn job descriptions. Their study mainly focused on collocated keywords and frequency counts using statistical approaches. While these methods are well established and documented, they may fail to capture higher-level semantic meaning in the text and struggle to answer more complex questions about the profession’s requirements and expectations.
George et al. (2023) used a qualitative method of interviewing industry experts, university researchers, and students to examine opportunities for prompt engineers, including job requirements and the characteristics of these emerging roles. Although this approach is appropriate for the research questions, they provide no transparency regarding their interview protocols, participant selection criteria, or analytical framework. This lack of transparency throughout the research process raises questions about the credibility of their results.
With the rise of LLMs, significantly more researchers have approved and applied these models to qualitative research (Morgan, 2023; Kirsten et al., 2025; Zhang et al., 2024). Fecher et al. (2025) empirically explored the implications of large language models on the science system, including the required competencies for their effective use. Several studies (e.g., Kirsten et al. (2025); Zhang et al. (2024)) provide their own frameworks and approaches to effectively apply LLMs to qualitative research methods. This paper takes a mix of some of the suggested approaches and adapts them to our specific needs.
@@ 2.2 Skill Extraction Techniques
Figure 1: Skill extraction results using SkillSpan (Zhang et al., 2022b). The highlighted text spans show extracted skills with contextual information that complicates downstream analysis.
Zhang et al. (2022b) applied an approach based on sequence labeling. Their model, SkillSpan, is a BERT-based model that is fine-tuned on manually annotated job postings. Their approach reached 57.7% F1 score, and their approach also includes transparent annotation guidelines (Zhang et al., 2022b) and was later evaluated by other researchers (e.g., Nguyen et al. (2024)). However, their approach faces a critical limitation: because of the way they annotated skills, the extracted skills often include long spans containing extraneous information. As shown in Figure 1, the identified skills—while being relevant—include contextual elements complicating downstream analysis such as skill categorization.
@@ 2.3 Skill Extraction with LLMs
Gnehm et al. (2022) presented an approach for extracting and classifying skill requirements in German-speaking job postings. Their method adopts pre-trained transformer-based language models for extracting skills, using context from job postings and the ESCO taxonomy le Vrang et al. (2014). As with Kompetencer (Zhang et al., 2022a), the approach is limited as it draws on an existing domain taxonomy to match potentially novel skills.
@@ 3.1 Data Collection
For this study, we use a dataset of job postings scraped from LinkedIn on April 7, 2025. The one-day crawl provides us a snapshot view into the online job market on this given date. While job postings on LinkedIn do not reflect all jobs advertised online, LinkedIn represents one of the largest global platforms for professional networking and job advertisements, making it a valuable source for analyzing current labor market trends. The data was crawled across the 60 countries with the most technological expertise, based on the U.S. News & World Report rankings (U.S. News, 2025). The countries are listed in Appendix A. The distribution of jobs reflects the concentration of tech jobs world-wide. For our comparative analysis, we selected four job roles for being highly relevant and potentially overlapping with the role of prompt engineer. These job roles are data engineer, data analyst, machine learning engineer, and data scientist.
The data collection was conducted in a 2 steps process: First we gathered all links to job postings that match the search criteria, then we extracted the detailed information for each job posting. For identifying job roles, we searched for exact matches in job titles, using quoted keywords to ensure precise matching jobs (e.g., “prompt engineer”). This approach allowed us to create a precise snapshot of job market for these five roles at a specific point in time.
@@ 3.2.1 Deduplication
Figure 2 displays the distribution of similarity scores across all job posting pairs. Based on visual inspection of this distribution, we can identify a distinct secondary mode at a threshold of 0.95 (highlighted with a dashed red line), above which job postings were considered duplicates.
Figure 2: Distribution of similarity scores between job posting pairs. The red vertical line indicates our chosen threshold (0.95) for identifying duplicate postings.
@@ 3.3 Skill Extraction
To address RQ2 (What skills and qualifications are required for the role?), we initially attempted to apply the SkillSpan model (Zhang et al., 2022b) for skill extraction. However, this approach proved suboptimal for our specific dataset. The model frequently extracted entire phrases as skill entities (see Figure 1), making it hard to identify the skill in our analysis. Combined with the relatively low accuracy of the model when applied to our dataset, SkillSpan was not a feasible option to extract the skills from the job postings.
@@ 3. Categorize each identified item into the appropriate category:
Figure 3: Prompt for skill extraction from job postings
Instead, we followed the qualitative analysis framework proposed by Zhang et al. (2024), which suggests specific techniques for designing effective LLM prompts for qualitative analysis of data. Figure 3 showcases our prompt, which has the following key elements:
@@
The extracted skills contained some variation in terminology and phrasing. For instance, “Python programming,” “Python development,” and “coding in Python” all refer to the same skill. To standardize the extracted information, we implemented an approach using k-means clustering to group semantically similar skills together. Each extracted skill was first mapped to a high-dimensional vector using sentence transformers (Reimers and Gurevych, 2019), specifically the “all-MiniLM-L6-v2” variant which builds upon the MiniLM architecture (Wang et al., 2020). This model creates embeddings that capture the semantics of text phrases, meaning that skills with similar meanings are located close to each other in the embedding space. We then applied k-means clustering to these skill vectors. K-means partitions the vector space into k clusters by minimizing the within-cluster variance, assigning each skill to the nearest cluster centroid. To determine an appropriate number of clusters, we conducted an iterative evaluation: we varied k from 5 to 20 and manually examined the top terms in each cluster to assess semantic coherence. Based on this evaluation, we selected 8 clusters for soft skills and 12 clusters for hard skills, as these provided the most interpretable groupings.
We visualize the embedding clusters with UMAP (McInnes et al., 2018), a dimensionality reduction technique that maps high-dimensional data (in our case, embeddings) to a two-dimensional space while preserving both local and global structures. UMAP was found to have advantages over techniques such as t-SNE and Principal Component Analysis (Coenen and Pearce, nd; McInnes et al., 2018).
@@ 4.1 Descriptive Data Overview (RQ1)
After removing duplicates, our final dataset had 20,662 unique job postings across five job roles (prompt engineer, machine learning engineer, data scientist, data engineer, and data analyst). Figure 4 shows how these jobs were distributed in the dataset. Despite the high relevance of LLMs in the tech industry today, prompt engineer positions make up a tiny part of the job market (72 jobs), compared to other data and AI roles, such as data engineer (8,272 jobs), data analyst (6,782 jobs), data scientist (5,524 jobs), and machine learning engineer (1,958 jobs).
Figure 4: Distribution of job postings across roles after removal of duplicates
@@
Figure 5: Skills cluster distribution for all analyzed jobs
This distribution confirms that while LLM use has grown quickly in industry, dedicated prompt engineer positions are still rare in today’s job market, making up less than 0.5% of related roles in our dataset. This matches our expectation that the role is still emerging and has not, yet, become as established as older positions.
@@ 4.2 Skills Analysis Across Job Roles
To understand what makes prompt engineering unique as a novel profession, we analyzed and compared both soft skills and hard skills required across different job roles. This analysis is contrasted with an analysis of the distinct skill profiles for prompt engineers (see Section 4.3), giving insights into how these roles differ despite working in related areas.
@@ 4.2.1 Soft skills requirements across jobs
Figure 5(a) and Table 2 show the results of our soft skill analysis across all collected jobs. Our k-means clustering identified eight distinct groups of soft skills. Statistical testing confirms significant differences between the skills profiles of the different job roles ($\chi^{2}(28,N=20662)=1377.78$, $p<0.0001$), validating that the skill requirements variation across roles is not due to random chance.
@@ 4.2.2 Hard skills requirements across jobs
7(b) and Table 4 depict the results of our comparative hard skill analysis across job postings. We identified twelve distinct hard skill clusters that appear consistently across job postings. Chi-square testing revealed highly significant differences among job types ($\chi^{2}(44,N=20662)=67291.55$, $p<0.0001$), confirming that the technical skill profiles across different job roles represent meaningful distinctions rather than random variation. This significant variation reflects the specialized technical requirements of each job position. Big Data Processing (25.1%) emerges as the most frequently required technical skill cluster across all job roles, followed by Data Warehousing & Modeling (12.6%) and SQL & Database Systems (8.7%). These three clusters form the foundation of technical requirements for data professionals, accounting for approximately 46% of all hard skills mentioned in job postings.
@@ 4.3.1 Soft skills requirements for prompt engineers (RQ2)
Figure 6: Comparative distribution of skills clusters across job roles
As depicted in 6(a), 7(a) and Table 2, prompt engineer roles place a particularly strong emphasis on collaborative communication, which together accounts for 43.8% of their soft-skill requirements—higher than the $\approx 39\%$ observed in other roles. This reflects how prompt engineers must work closely with stakeholders to translate practical needs into effective language-model prompts.
@@ 4.3.2 Hard skills requirements for prompt engineers (RQ3)
Figure 7: Skills cluster distribution for prompt engineer positions, in comparison to skills in other jobs
Figure 6(b) and tables 4 and 4 present a side-by-side comparison of hard skill requirements across all five job roles, further illustrating the distinctive technical profile of prompt engineers. This comparison reveals several key insights:
  • Prompt engineers require the highest proportion of Machine Learning & AI skills (22.8%) compared to machine learning engineers (21.6%) and data scientists (16.0%), yet their focus is specifically on language models rather than general machine learning algorithms.
@@
The emergence of these specialized skill requirements confirms that prompt engineering is developing as a distinct technical discipline rather than simply being an extension of existing job roles. This technical specialization, combining AI expertise with prompt-specific methods and tools, supports our finding that prompt engineering represents a new professional category with its own distinct skill profile, even as it remains a relatively small segment of the current job market.

### v2

# Prompt Engineer: Analyzing Hard and Soft Skill Requirements in the AI Job Market

@@ Abstract
The rise of large language models (LLMs) has created a new job role: the Prompt Engineer. Despite growing interest in this position, we still do not fully understand what skills this new job role requires or how common these jobs are. In this paper, we present a data-driven analysis of global prompt engineering job trends on LinkedIn. We take a snapshot of the evolving AI workforce by analyzing 20,662 job postings on LinkedIn, including 72 prompt engineer positions, to learn more about this emerging role. We find that prompt engineering is still rare (less than 0.5% of sampled job postings) but has a unique skill profile. Prompt engineers need AI knowledge (22.8%), prompt design skills (18.7%), good communication (21.9%), and creative problem-solving (15.8%) skills. These requirements significantly differ from those of established roles, such as data scientists and machine learning engineers. Our findings help job seekers, employers, and educational institutions in better understanding the emerging field of prompt engineering.
Keywords: AI workforce; human-AI interaction; job market analysis; large language models; LLM; natural language processing; prompt engineering; skill analysis; text mining.
@@ 1 Introduction
The rise of large language models (LLMs) has been transforming various industries (Chkirbene et al., 2024). AI enables advancements in automation, content generation, and AI-driven decision-making. However, the effectiveness of AI models heavily depends on how prompts are structured (Oppenlaender, 2023), leading to the emergence of prompt engineering as a specialized field. Prompts are programs that steer the pre-trained LLM (Karpathy, 2025). Some companies are seeking to recruit prompt engineers in an effort to optimize and extract the full potential of LLMs. Despite the growing demand for this role, there remains no standardized definition of a prompt engineer, resulting in uncertainties regarding the scope, required skills, and industry relevance of this emerging job role.
Due to the timely relevance of LLM applications in industry, it is important to develop an understanding of the required knowledge and skills for this emerging job type. However, because of the novelty of AI technology, individuals who are interested in pursuing the role of the prompt engineer may face challenges in understanding the qualifications required. Therefore, addressing this gap is beneficial to both job seekers and practitioners in industry as well as employers who want to identify suitable candidates. However, how can we systematically extract and analyze the skills, competencies, and industry relevance of this job role? In this paper, we employ a text mining approach to analyze job advertisements published on LinkedIn, aiming to uncover the key knowledge areas and skills required for prompt engineers. Understanding these trends provides insight into the broader adoption of LLMs and the pace of different industries in adapting to this evolving technology.
@@
  • RQ1: How prevalent are prompt engineer positions compared to more established job roles?
  • RQ2: What soft skills are most valued for prompt engineers, and how do these differ from other job roles?
  • RQ3: What technical skills define the prompt engineer role, and how do these hard skills create a distinct professional profile?
@@
The remainder of this paper is structured as follows. We review related literature on prompt engineering and the extraction of skills in Section 2. We then describe our approach to collecting and extracting information from job descriptions in Section 3. We present our findings in Section 4, contrasting the skills of prompt engineers with five other related jobs. We discuss our findings in Section 5 and conclude in Section 6.
@@ 2.1 The Emergence of the Prompt Engineer
Artificial intelligence (AI) is changing the way workers perform their jobs and the skills they require (OECD, 2024; Brasse et al., 2024). Generative artificial intelligence has a strong impact on the future of work (Cazzaniga, 2024). While concerns and anxiety over automation are growing (Dinnen, 2022), AI also creates new professional opportunities (Kutela et al., 2023b). There is a demand for a new digital workforce with a novel set of AI literacy (Long and Magerko, 2020) and digital skills (van Laar et al., 2017). One such novel job is the prompt engineer.
@@
Kutela et al. (2023a) examined the nature, qualifications, and compensation of prompt engineers through an analysis of LinkedIn job descriptions. Their study mainly focused on collocated keywords and frequency counts using statistical approaches. While these methods are well established and documented, they may fail to capture higher-level semantic meaning in the text and struggle to answer more complex questions about the profession’s requirements and expectations.
George et al. (2023) used a qualitative method of interviewing industry experts, university researchers, and students to examine opportunities for prompt engineers, including job requirements and the characteristics of these emerging roles. Although this approach is appropriate for the research questions, they provide no transparency regarding their interview protocols, participant selection criteria, or analytical framework, making their findings hard to replicate.
With the rise of LLMs, significantly more researchers have approved and applied LLMs to qualitative research (Morgan, 2023; Kirsten et al., 2025; Zhang et al., 2023). Several studies (e.g., Kirsten et al. (2025); Zhang et al. (2023)) provide their own frameworks and approaches to effectively apply LLMs to qualitative research methods. This paper takes a mix of some of the suggested approaches and adapts them to our specific needs.
@@ 2.2 Skill Extraction Techniques
Zhang et al. (2022b) applied an approach based on sequence labeling. Their model, SkillSpan, is a BERT-based model that is fine-tuned on manually annotated job postings. Their approach reached 57.7% F1 score, and their approach also includes transparent annotation guidelines (Zhang et al., 2022b) and was later evaluated by other researchers (e.g., Nguyen et al. (2024)). However, their approach faces a critical limitation: because of the way they annotated skills, the extracted skills often include long spans of text containing extraneous information. However, the identified skills, while being relevant, include contextual elements complicating downstream analysis such as skill categorization.
@@ 2.3 Skill Extraction with LLMs
Gnehm et al. (2022) presented an approach for extracting and classifying skill requirements in German-speaking job postings. Their method adopts pre-trained transformer-based language models for extracting skills, using context from job postings and the ESCO taxonomy (le Vrang et al., 2014). As with Kompetencer (Zhang et al., 2022a), the approach is limited as it draws on an existing domain taxonomy to match potentially novel skills.
@@ 3.1 Data Collection
For this study, we use a dataset of job postings scraped from LinkedIn on April 7, 2025. The one-day crawl provides us a snapshot view into the online job market on this given date. While job postings on LinkedIn do not reflect all jobs advertised online, LinkedIn represents one of the largest global platforms for professional networking and job advertisements, making it a valuable source for analyzing current labor market trends. The data was crawled across the 60 countries with the most technological expertise, based on the U.S. News & World Report rankings (U.S. News, 2025). The countries are listed in the supplemental material. The distribution of jobs reflects the concentration of tech jobs world-wide. For our comparative analysis, we selected four job roles for being highly relevant and potentially overlapping with the role of prompt engineer. These job roles are data engineer, data analyst, machine learning engineer, and data scientist.
The data collection was conducted in a two-step process: First we gathered all links to job postings that match the search criteria, then we extracted detailed information for each job posting. For identifying job roles, we searched for exact matches in job titles, using quoted keywords to ensure precise matching jobs (e.g., “prompt engineer”). This approach allowed us to create a precise snapshot of job market for these five roles at a specific point in time.
@@ 3.2.1 Deduplication
Based on visual inspection of the distribution of similarity scores across all job posting pairs, we can identify a distinct secondary mode at a threshold of 0.95, above which job postings were considered duplicates.
@@ 3.3 Skill Extraction
To address RQ2 (What skills and qualifications are required for the role?), we initially attempted to apply the SkillSpan model (Zhang et al., 2022b) for skill extraction. However, this approach proved suboptimal for our specific dataset. The model frequently extracted entire phrases as skill entities, making it hard to identify the skill in our analysis. Combined with the relatively low accuracy of the model when applied to our dataset, SkillSpan was not a feasible option to extract the skills from the job postings.
@@ 3. Categorize each identified item into the appropriate category:
Figure 1: Prompt for skill extraction from job postings. Source: Authors own work.
Instead, we followed the qualitative analysis framework proposed by Zhang et al. (2023). This framework suggests specific techniques for designing effective LLM prompts for qualitative analysis of data. Figure 1 showcases our prompt, which has the following key elements:
@@
The extracted skills contained some variation in terminology and phrasing. For instance, “Python programming,” “Python development,” and “coding in Python” all refer to the same skill. To standardize the extracted information, we implemented an approach using k-means clustering to group semantically similar skills together. Each extracted skill was first mapped to a high-dimensional vector using sentence transformers (Reimers and Gurevych, 2019), specifically the all-MiniLM-L6-v2 variant which builds upon the MiniLM architecture (Wang et al., 2020). This model creates embeddings that capture the semantics of text phrases, meaning that skills with similar meanings are located close to each other in the embedding space. We then applied k-means clustering to these skill vectors. K-means partitions the vector space into k clusters by minimizing the within-cluster variance, assigning each skill to the nearest cluster centroid. To determine an appropriate number of clusters, we conducted an iterative evaluation: we varied k from 5 to 20 and manually examined the top terms in each cluster to assess semantic coherence. Based on this evaluation, we selected 8 clusters for soft skills and 12 clusters for hard skills, as these provided the most interpretable groupings.
We visualize the embedding clusters with UMAP (McInnes et al., 2018), a dimensionality reduction technique that maps high-dimensional data (in our case, embeddings) to a two-dimensional space while preserving both local and global structures. UMAP was found to have advantages over techniques such as t-SNE and Principal Component Analysis (Coenen and Pearce, n.d.; McInnes et al., 2018).
@@ 4.1 Descriptive Data Overview (RQ1)
After removing duplicates, our final dataset had 20,662 unique job postings across five job roles (prompt engineer, machine learning engineer, data scientist, data engineer, and data analyst). Despite the high relevance of LLMs in the tech industry today, prompt engineer positions make up a tiny part of the job market (72 jobs), compared to other data and AI roles, such as data engineer (8,272 jobs), data analyst (6,782 jobs), data scientist (5,524 jobs), and machine learning engineer (1,958 jobs).
@@
Figure 2: Skills cluster distribution for all analyzed jobs.
This distribution confirms that while LLM use has grown quickly in industry, dedicated prompt engineer positions are still rare in today’s job market, making up less than 0.5% of related roles in our dataset. This matches our expectation that the emerging role of ‘Prompt Engineer’ has not become as established as older positions.
@@ 4.2 Skills Analysis Across Job Roles
To understand what makes prompt engineering unique as a potential novel profession, we analyzed and compared both soft skills and hard skills required across different job roles. This analysis is contrasted with an analysis of the distinct skill profiles for prompt engineers (see Section 4.3), giving insights into how these roles differ despite working in related areas.
@@ 4.2.1 Soft skills requirements across jobs
Figure 2a and Table 2 show the results of our soft skill analysis across all collected jobs. Our k-means clustering identified eight distinct groups of soft skills. Statistical testing confirms significant differences between the skills profiles of the different job roles ($\chi^{2}(28,N=20662)=1377.78$, $p<0.0001$), validating that the skill requirements variation across roles is not due to random chance.
@@ 4.2.2 Hard skills requirements across jobs
Figure 4b and Figure 4 depict the results of our comparative hard skill analysis across job postings. We identified twelve distinct hard skill clusters that appear consistently across job postings. Chi-square testing revealed highly significant differences among job types ($\chi^{2}(44,N=20662)=67291.55$, $p<0.0001$), confirming that the technical skill profiles across different job roles represent meaningful distinctions rather than random variation. This significant variation reflects the specialized technical requirements of each job position. Big Data Processing (25.1%) emerges as the most frequently required technical skill cluster across all job roles, followed by Data Warehousing & Modeling (12.6%) and SQL & Database Systems (8.7%). These three clusters form the foundation of technical requirements for data professionals, accounting for approximately 46% of all hard skills mentioned in job postings.
@@ 4.3.1 Soft skills requirements for prompt engineers (RQ2)
Figure 3: Comparative distribution of skills clusters across job roles. Source: Authors own work.
As depicted in Figure 3a, Figure 4a and Figure 2, prompt engineer roles place a particularly strong emphasis on collaborative communication, which together accounts for 43.8% of their soft-skill requirements—higher than the $\approx 39\%$ observed in other roles. This reflects how prompt engineers must work closely with stakeholders to translate practical needs into effective language-model prompts.
@@ 4.3.2 Hard skills requirements for prompt engineers (RQ3)
Figure 4: Skills cluster distribution for prompt engineer positions, in comparison to skills in other jobs. Source: Authors own work.
Figure 3b and tables 4 and 4 present a side-by-side comparison of hard skill requirements across all five job roles, further illustrating the distinctive technical profile of prompt engineers. This comparison reveals several key insights:
  • Prompt engineers require a high proportion of Machine Learning & AI skills (22.8%), even higher than machine learning engineers (21.6%) and data scientists (16.0%), yet their focus is specifically on language models rather than general machine learning algorithms.
@@
The emergence of these specialized skill requirements confirms that prompt engineering has a distinct technical profile rather than simply being an extension of existing job roles. This technical specialization, combining AI expertise with prompt-specific methods and tools, supports our finding that prompt engineering is a representative of a new professional category with its own distinct skill profile, even as it remains a relatively small segment of the current job market.